### v1

# 3-manifolds with universal links are spherical

Abstract.
We determine which orientable $3$-manifolds are dominated by the $3$-sphere. This characterization implies that the closed $3$-manifolds that admit a universal link are spherical. Moreover, we show examples of complement universal links that are not universal.
1. Introduction
Hilden and Montesinos proved that every orientable $3$-manifold is a $3$-fold branched covering of the $3$-sphere $\mathbb{S}^{3}$. With this result as a starting point, Thurston proved the striking fact that there exists a link $L\subset\mathbb{S}^{3}$ such that every orientable $3$-manifold is a branched covering of $\mathbb{S}^{3}$ with branching set exactly $L$. Thurston called $L$ a universal link. In this paper, we characterize the existence of universal links for closed $3$-manifolds:
Corollary 4.2.
A closed $3$-manifold admits a universal link if and only if it is spherical.
This result is a consequence of a more general result in non-zero degree maps, where branched covers between orientable closed manifolds are a particular case. We define a reflexive and transitive relation between orientable closed manifolds by saying that $N$ dominates $M$ if there is a non-zero degree map from $N$ to $M$ (see Section 4). The existence of universal links implies that all orientable $3$-manifolds dominate the $3$-sphere. By definition, all spherical $3$-manifolds are dominated by the $3$-sphere, but actually these are the only $3$-manifolds that satisfy this property:
Theorem 4.1.
The top elements in the set of closed, orientable $3$-manifolds, under the domination relation, are exactly the spherical $3$-manifolds.
We give two proofs of Theorem 4.1, one using the residual finiteness for compact $3$-manifolds [Hempel:ResidualFnite], and a second using [KotschickNeofytidis:DominationByCircleBunles, Theorem 3] due to Kotschick and Neofytidis, which allows us to construct a finite cover from a connected sum of $\mathbb{S}^{2}\times\mathbb{S}^{1}$ to manifolds without aspherical prime factors in their prime decomposition.
It is worth mentioning that Edmonds proved in [Edmonds:Pi_1SurjectiveMapsBranchedCovers] that every map between closed, orientable 3-manifolds with a degree greater than $2$ is homotopic to a branched cover, which explains a reciprocal relation between non-zero degree maps and branched covers.
Finally, we notice that Thurston provided a more general definition of a universal link in [Thurston:UniversalLinks], by calling a link $L$ in an arbitrary orientable closed $3$-manifold $M$ complement universal if there exists a link in any orientable closed $3$-manifold whose complement is virtually homeomorphic to $M{\smallsetminus}L$ (see Section 5). We indeed showed that these two definitions are not equivalent:
Corollary 5.3.
There exists a universal link in infinitely many $3$-manifolds which is not a complement universal link.
1.1. Acknowledgements
We thank Mario Eudave for pointing us toward helpful references.
2. Universal links in spherical $3$-manifolds
We begin by defining branched coverings between 3-manifolds and universality.
Definition 2.1.
An open proper map $\phi:M^{3}\to N^{3}$ is called a branched covering if there exists a link $L\subset N^{3}$ such that the restriction $\phi|:M^{3}-\varphi^{-1}(L)\to N^{3}-L$ is a finite covering space. The link $L$ is called the branching set.
Definition 2.2.
We define a link $L$ in a closed $3$-manifold $M$ to be universal if every orientable closed $3$-manifold NN is a branched cover of $M$ over $L.$
Definition 2.3.
A spherical $3$-manifold is the resultant quotient space $\mathbb{S}^{3}/\Gamma$, where $\Gamma$ is a finite subgroup of $SO(4)$ acting freely by rotations on $\mathbb{S}^{3}$. Special cases of these manifolds include lens spaces and prism manifolds.
Theorem 2.4.
Let $M$ be a spherical 3-manifold; then $M$ has universal knots and links.
Proof.
We address the proof of the existence of universal links; however, the same can be applied to knots.
Let $M=\mathbb{S}^{3}/\Gamma$ be a spherical manifold, where $\Gamma$ is a finite subgroup of $SO(4)$ acting freely on $\mathbb{S}^{3}$. Thus, the projection map $\pi\colon\mathbb{S}^{3}\to\mathbb{S}^{3}/\Gamma$ is a covering map.
Let $k\subset\mathbb{S}^{3}$ be a universal link. We can find an isotopy of $\mathbb{S}^{3}$ such that the image of $k$ under this isotopy, say $K$, is contained in an open ball $B$, and $\pi|_{B}\colon B\to\pi(B)$ is a homeomorphism. In this way, $\pi(K)=K^{\prime}$ is a link in $M=\mathbb{S}^{3}/\Gamma$. We state that $K^{\prime}$ is a universal link in $M$:
[…]
In the following section, we explore the obstruction of the existence of branched covering maps between the $3$-sphere and an orientable $3$-manifold in a wider family of maps, which are the ones with non-zero degree.
3. Transfer maps
The subject of transfer maps on covering maps is well-known. But its existence can be extended to non-zero degree maps. The results presented in this section are probably folklore knowledge among experts. However, it might not be easy to extract from the literature. Although we could track the first moment this argument was used in the proof of [BeresteinEdmonds:BranchedCovers, Theorem 7.3]. For this reason, and for the sake of being more self-contained, we provide the details here.
Definition 3.1.
Given a map $f:M^{n}\to N^{n}$ between orientable $n$-manifolds, we define the transfer of
$f_{\#}:H_{k}(M)\to H_{k}(N)$
as the map
$\tau_{f}=PD_{M}\circ f^{\#}\circ PD_{N}^{-1}:H_{k}(N)\to H_{k}(M),$
where $PD_{*}$ denotes the Poincaré duality isomorphism, and $f_{\#},f^{\#}$ are the induced maps of $f$ in homology and cohomology, respectively.
Notice that $\tau_{f}$ is the map from $H_{k}(N)$ to $H_{k}(M)$ obtained by following the commutative diagram below, without passing through $f_{\#}$ directly:
$H_{k}(M)$$H_{k}(N)$$H^{n-k}(M)$$H^{n-k}(N)$$\scriptstyle{f_{\#}}$$\scriptstyle{PD_{M}}$$\scriptstyle{f^{\#}}$$\scriptstyle{PD_{N}}$
We now prove the following result, where $\deg(f)$ denotes the degree of $f$. Recall that this degree is defined for any map between orientable manifolds $f:M^{n}\to N^{n}$ as the integer $d$ satisfying
$f_{\#}([M])=d\cdot[N]\in H_{n}(N),$
where $M]\in H_{n}(M)$ and $N]\in H_{n}(N)$ are the fundamental classes of $M$ and $N$, respectively.
Lemma 3.2.
$f_{\#}\circ\tau_{f}([\alpha])=\deg(f)\cdot[\alpha]$.
Proof.
This follows directly from the definition $\tau_{f}:=PD_{M}\circ f^{\#}\circ PD_{N}^{-1}$.
Since $PD_{N}$ is an isomorphism, there exists a unique class $a]\in H^{n-k}(N)$ such that
$PD_{N}([a])=[\alpha]\quad\text{i.e.,}\quad[a]\smallfrown[N]=[\alpha].$
Then we compute:
$f_{\#}\circ\tau_{f}([\alpha])$ $=f_{\#}\circ\tau_{f}([a]\smallfrown[N])$
$=f_{\#}\circ(PD_{M}\circ f^{\#}\circ PD^{-1}_{N})([a]\smallfrown[N])$
$=f_{\#}\circ(f^{\#}([a])\smallfrown[M])$
By the naturality of the cap product, $f_{\#}(f^{\#}([a])\smallfrown[M])=[a]\smallfrown f_{\#}([M])$, and by definition, $f_{\#}([M])=\deg(f)\cdot[N]$. Therefore,
$f_{\#}\circ\tau_{f}([\alpha])$ $=([a]\smallfrown f_{\#}([M]))$
$=([a]\smallfrown deg(f)\cdot[N])$
$=\deg(f)\cdot([a]\smallfrown[N])$
$=\deg(f)\cdot[\alpha]$
Observe that if $f:M\to N$ is an $d$-sheeted branched covering map, then $\deg(f)=d$. In particular, when $M=\mathbb{S}^{3}$, this implies that $f_{\#}\circ\tau_{f}:H_{1}(N)\to H_{1}(N)$ is simultaneously the zero map and the multiplication by $d$. This yields the following result.
Lemma 3.3.
Let $N$ be a orientable $3$-manifold. If there exists a map $f:\mathbb{S}^{3}\to N$ of non-zero degree, then $H_{1}(N)$ is a finite group; equivalently, $N$ is a rational homology sphere.
Lemma 3.4.
Let $\phi:\widetilde{M}\to M$ be a finite covering map, and let $f:\mathbb{S}^{3}\to M$ be a map of non-zero degree. Then there exists a lift $\widetilde{f}:\mathbb{S}^{3}\to\widetilde{M}$ of $f$ with non-zero degree.
[…]
Proof.
In homology, we have the relation $\phi_{\#}\circ\widetilde{f}_{\#}=f_{\#}$. Taking degrees gives
$\deg(\phi)\cdot\deg(\widetilde{f})=\deg(f)\neq 0,$
so $\deg(\widetilde{f})\neq 0$. ∎
In [Smith:TransferRamifiedCoverings], L. Smith defined an $d$-fold ramified covering map, where branched coverings are a particular case, and proved a generalization of Lemma 3.2 that includes non-orientable manifolds, as follows:
Proposition 3.5 (Proposition 2.2 in [Smith:TransferRamifiedCoverings]).
Let $f:M^{n}\to N^{n}$ be a $d$-fold ramified covering between $n$-manifolds, then there exist a transfer homomorphism of
$f_{\#}:H_{k}(M)\to H_{k}(N)$
denoted as
$\tau_{f}:H_{k}(N)\to H_{k}(M),$
such that
$f_{\#}\circ\tau_{f}:H_{k}(N)\to H_{k}(N),$
is multiplication by $d$.
4. Top elements of closed $3$-manifolds under domination relation
The existence of a non-zero degree map, $N\rightarrow M$, defines a reflexive and transitive relation on the set of homotopy types of closed orientable manifolds. Whenever such a map exists, we say that $N$ dominates $M$ and write $N\geq M$. In this case $N$ is at least as complicated as $M$. For example, in Lemma 3.2 we saw that the induced maps in rational homology are surjective, thus, in particular, the Betti numbers of $M$ are bounded from above by those of $N$. Also, $N\geq M$ implies that the fundamental group of $N$ surjects onto a finite index subgroup of the fundamental group of $M$.
Theorem 4.1.
The top elements in the set of closed, orientable $3$-manifolds, under the domination relation, are exactly the spherical $3$-manifolds.
As every $3$-manifold is dominated by spherical $3$-manifolds by Theorem 2.4, it is enough to show that if $M$ is dominated by $\mathbb{S}^{3}$ then $M$ has a finite fundamental group. We provide two proofs of this fact:
First proof.
As non-spherical irreducible, closed, orientable $3$-manifolds are aspherical, then there is no domination of $\mathbb{S}^{3}$ to them. So we focus on the reducible case.
On one hand, if $M$ is prime, then $M=\mathbb{S}^{1}\times\mathbb{S}^{2}$ and it can not be dominated by $\mathbb{S}^{3},$ because of Lemma 3.3.
On the other hand, if $M$ has a separating $2$-sphere $S,$ such that $M=M_{1}\sharp_{S}M_{2}$, the residual finiteness for compact $3$-manifolds [Hempel:ResidualFnite], implies that there is $K_{i}$ a proper normal subgroup of $\pi_{1}(M_{i})$ of finite index, for each $i=1,2$.
Let $\phi:N\rightarrow M$ be the finite cover induced by the kernel of the following epimorphism:
$\pi_{1}(M_{1})\ast\pi_{1}(M_{2})\rightarrow(\pi_{1}(M_{1})/K_{1})\times(\pi_{1}(M_{2})/K_{2})$
Consider $p\in S,$ $q\in\phi^{-1}(p)$ and $S_{1}$ the connected component of $\phi^{-1}(S)$ that contains $q.$
We claim that that $S_{1}$ is a non-separating $2$-sphere. Indeed, take a simple loop $x_{i}:([0,1],0)\rightarrow(M_{i},p)$ whose homotopy class belongs to $\pi_{1}(M_{i}){\smallsetminus}K_{i}.$ Notice that the homotopy class $x_{1}][x_{2}][x_{1}]^{-1}[x_{2}]^{-1}\in Ker(\phi).$ Then the corresponding loop, lifts to a loop in $N,$ denoted by $s.$ Notice that neither $x_{1}],[x_{1}][x_{2}]$ nor $x_{1}][x_{2}][x_{1}]^{-1}$ are in $Ker(\phi),$ so $s$ intersects once $S_{1}$ at $q.$ And the intersection is transversal, because $s$ starts inside $\phi^{-1}(M_{1})$ and ends inside $\phi^{-1}(M_{2}),$ which are in different sides of $S_{1}.$
After taking a regular neighborhood of $s\cup S_{1},$ we have that NN has $\mathbb{S}^{1}\times\mathbb{S}^{2}$ as a prime summands in its prime decomposition, but by Lemma 3.4, and Lemma 3.3 we have that $N$ is a rational homology sphere, which is impossible. ∎
Second proof.
By the Kneser-Milnor prime decomposition, $M$ has an essentially unique prime decomposition $M=M_{1}\sharp...\sharp M_{k}$ under the connected sum operation. Each prime summand $M_{i}$ is either aspherical, $\mathbb{S}^{1}\times\mathbb{S}^{2},$ or has a finite fundamental group.
Suppose we have a non-zero degree map $f:\mathbb{S}^{3}\rightarrow M.$ If $M$ has an aspherical summand $M_{i}$ in its prime decomposition, then by composing $f$ with the degree one map $M\rightarrow M_{i}$ that collapses the connected summands other than $M_{i},$ we obtain a dominant map $\mathbb{S}^{3}\rightarrow M_{i},$ but that is impossible because $\pi_{3}(M_{i})=\{0\}$.
So, if we have a non-zero degree map $f:\mathbb{S}^{3}\rightarrow M,$ then $M$ does not have an aspherical summand $M_{i}$ in its prime decomposition. By [KotschickNeofytidis:DominationByCircleBunles, Theorem 3] we conclude that $M$ is finitely covered by a connected sum $\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ where the empty connected sum $(k=0)$ denotes the $3$-sphere $\mathbb{S}^{3}.$
Notice that if $k>0$ then we could lift $f:\mathbb{S}^{3}\rightarrow M$ to a non-zero degree map $\widetilde{f}:\mathbb{S}^{3}\rightarrow\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ by Lemma 3.4, and Lemma 3.3 would imply that $\sharp_{k}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ is a rational homology sphere, which is impossible. Then $k=0$, meaning that $M$ is a spherical 33-manifold. ∎
Corollary 4.2.
A closed $3$-manifold admits universal links if and only if it is spherical.
Proof.
By Theorem 4.1 and Theorem 2.4 we have the case when $M$ is an orientable closed 33-manifold. For the non-orientable case, notice that $M$ non-zero first Betti number [BeresteinEdmonds:BranchedCovers, Lemma 8.6]. Then by Proposition 3.5, we have that there is not branching cover from the $3$-sphere to $M.$
5. Complement universal links that are not universal
In [Thurston:UniversalLinks] Thurston gave a more general definition of a universal link, as follows:
Definition 5.1.
We define a link $L$ in a orientable closed $3$-manifold $M$ to be complement universal if every orientable closed $3$-manifold NN contains a link $L^{\prime}\subset N$ such that $N{\smallsetminus}L^{\prime}$ is homeomorphic to a finite sheeted covering space of $M{\smallsetminus}L.$
Remark 5.2.
To avoid confusion on the concepts. Our concept of universal, which is widely used in literature on branching covers, is called in [Thurston:UniversalLinks] under the name of universal branching and our concept of complement universal is called in the same paper under the name of universal.
Notice that if $L$ is a universal link, it is a complement universal link, because the restriction of a branched covering map on the complement of the branching set is a finite covering map (see Def. 2.2).
Corollary 5.3.
There exists a complement universal link in infinitely many $3$-manifolds that is not a universal.
Proof.
Consider $M$, a non-spherical $3$-manifold that is homeomorphic to a Dehn surgery over a hyperbolic universal link $L$ in a spherical $3$-manifold $S.$ By Thurston’s hyperbolic Dehn filling theorem, this provides infinitely many $3$-manifolds $M.$ Moreover, for every $N$ orientable closed $3$-manifold there is a branched cover $p:(N,p^{-1}(L))\rightarrow(S,L),$ which induces a finite cover from $N{\smallsetminus}p^{-1}(L)$ to $S{\smallsetminus}L,$ and the latter is homeomorphic to the complement of the core $L^{\prime}$ of the Dehn filled tori in $M.$
In conclusion, $L^{\prime}\subset M$ is a complement universal link, that is not a universal link because $M$ is non-spherical by Corollary 4.2. And by varying over the integral Dehn surgery on $L$ in $S$, we obtain infinitely many non-homeomorphic $3$-manifolds $M.$
6. Further comments
Berestein and Edmons proved in [BeresteinEdmonds:BranchedCovers], among other things, that an orientable $3$-manifold $M$ is a $3$-fold branched covering of $\mathbb{S}^{2}\times\mathbb{S}^{1}$ if and only if $H^{1}(M)\neq 0$. Some time later, Núñez in [Nuñez:UniversalLinksInS2XS1] gave an example of a link $L$ in $\mathbb{S}^{2}\times\mathbb{S}^{1}$ such that every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched covering over $\mathbb{S}^{2}\times\mathbb{S}^{1}$ with branching set exactly the link $L,$ and we denote these links as universal link for $\mathbb{S}^{2}\times\mathbb{S}^{1}.$ In the case of aspherical $3$-manifolds we ask:
Question 6.1.
Does there exist a closed, aspherical, orientable $3$-manifold $M$ and a link $L$ inside, such that every closed, aspherical, orientable $3$-manifold is a branched covering over $M$ with branching set exactly the link $L$?
And the next problem:
Problem 6.2.
Determine all the closed orientable $3$-manifolds $M$ with a link $L$ inside such that every closed, orientable $3$-manifold NN with $H^{1}(N)\neq 0$ is a branched covering over $M$ with branching set exactly the link $L.$
As we saw in Theorem 5.3, the notions of universal link and universal branching link are not equivalent for non-spherical $3$-manifolds, but one could ask whether these notions coincide in the case of spherical $3$-manifolds.
Question 6.3.
Are there complement universal links in spherical $3$-manifolds that are not universal?
In the proof of Theorem 2.4, we construct universal knots in spherical 3-manifolds by taking a universal knot in $\mathbb{S}^{3}$ and embedding it into the manifold. This naturally raises the question of whether all universal knots in spherical manifolds arise in this way. More precisely, we can ask:
Question 6.4.
Do there exist universal knots in spherical 3-manifolds that are not contained in a 3-ball?